# Joint Self-Supervised Video Alignment and Action Segmentation

### v2

@@ Abstract
We introduce a novel approach for simultaneous self-supervised video alignment and action segmentation based on a unified optimal transport framework. In particular, we first tackle self-supervised video alignment by developing a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which trains efficiently on GPUs and needs only a few iterations for solving the optimal transport problem. Our single-task method achieves the state-of-the-art performance on multiple video alignment benchmarks and outperforms VAVA, which relies on a traditional Kantorovich optimal transport formulation with an optimality prior. Furthermore, we extend our approach by proposing a unified optimal transport framework for joint self-supervised video alignment and action segmentation, which requires training and storing a single model and saves both time and memory consumption as compared to two different single-task models. Extensive evaluations on several video alignment and action segmentation datasets demonstrate that our multi-task method achieves comparable video alignment yet superior action segmentation results over previous methods in video alignment and action segmentation respectively. Finally, to the best of our knowledge, this is the first work to unify video alignment and action segmentation into a single model.
@@ 1 Introduction
Though the past decade has witnessed remarkable progress in human activity understanding in videos, the majority of the research efforts have been invested in action recognition [8, 69, 76, 21], which categorizes simple actions in short videos. In this paper, we study the two less-explored problems, i.e., temporal video alignment (frame-to-frame assignment), which establishes framewise correspondences between long videos recording a complex activity, and temporal action segmentation (frame-to-action assignment), which assigns frames of long videos capturing a multi-phase activity to phase/action labels. Since acquiring per-frame annotations for supervised training is generally difficult and costly, we are interested in self-supervised approaches for video alignment and action segmentation.
@@
One popular group of self-supervised video alignment methods rely on global alignment techniques widely used in time series literature. For example, LAV [26] utilizes dynamic time warping [12] by assuming monotonic orderings and no background/redundant frames. VAVA [45] relaxes the above assumptions by incorporating an optimality prior into a standard Kantorovich optimal transport framework [11], along with an inter-video contrastive term and an intra-video contrastive term. However, it is challenging to balance multiple losses as well as handle repeated actions. Similarly, self-supervised action segmentation methods based on optimal transport have been introduced, including TOT [35] and UFSA [71] which may suffer in cases of order variations, unbalanced segmentation, and repeated actions. ASOT [80] addresses these drawbacks via a fused Gromov-Wasserstein optimal transport framework with a structural prior, which outperforms previous works in self-supervised action segmentation. Lastly, though both self-supervised video align and action segmentation require fine-grained temporal understanding of videos, their interaction in a multi-task learning setup has not been explored.
@@
Motivated by the above observations, we first propose VAOT (see Fig. 1(a)), a novel self-supervised video alignment approach based on a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which tackles order variations, background/redundant frames, and repeated actions in a single global alignment framework. Our single-task model trains efficiently on GPUs and requires few iterations to derive the optimal transport solution, while outperforming previous methods, including VAVA [45], on video alignment datasets. Moreover, we develop VASOT (see Fig. 1(b)), a joint self-supervised video alignment and action segmentation approach by exploring the relationship between self-supervised video alignment and action segmentation via a unified optimal transport framework. Our multi-task model performs on par with prior works on video alignment benchmarks yet establishes the new state of the art on action segmentation benchmarks. In addition, our joint model requires training and storing a single model and saves both time and memory consumption as compared to two separate single-task models. Lastly, we observe in Sec. 4.3 that, in a multi-task learning setting, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance significantly.
@@ 2 Related Work
Self-Supervised Learning. Early self-supervised learning methods focus on designing image-based pretext tasks with pseudo-labels as supervision signals for learning representations such as image colorization [37, 38], object counting [50, 46], predicting rotations [24], solving puzzles [5, 31], image inpainting [28], and image clustering [6, 7]. The above image-based approaches mostly extract spatial cues from the image content. Recently, significant efforts have been invested in video-based self-supervised learning methods, which exploit both spatial and temporal information. Examples of video-based pretext tasks include forecasting future frames [61, 74, 1, 15], enforcing temporal coherence [49, 84, 25], predicting temporal order [48, 39, 23, 78], arrow of time [53, 77], and pace [4, 75, 81], and utilizing contrastive learning [22, 27, 54, 13]. More recently, skeleton-based self-supervised learning methods with skeleton-based pretext tasks, e.g., skeleton inpainting [83], motion prediction [63], skeleton sequence alignment [36, 70], and utilizing neighborhood consistency [59], motion continuity [64], and multiple pretext tasks [44], have been introduced. These skeleton-based approaches may suffer from human pose estimation errors and missing context details. Here, we leverage video alignment and/or action segmentation as our video-based pretext tasks.
@@
Video Alignment. Self-supervised video alignment has attracted a great amount of research interest in recent years. TCC [19] enforces cycle consistencies between corresponding frames across videos for learning representations. Recently, GTCC [17] extends TCC [19] by proposing multi-cycle consistencies for tackling repeated actions. Both TCC [19] and GTCC [17] perform local alignment by aligning each frame separately. Motivated by global alignment techniques for time series, methods which align the video as a whole have been introduced. LAV [26] which assumes monotonic orderings and no background/redundant frames relies on dynamic time warping [12]. To handle non-monotonic orderings and background/redundant frames, VAVA [45] employs a traditional Kantorovich optimal transport formulation [11] with an optimality prior. In this work, we propose a fused Gromov-Wasserstein optimal transport formulation with a structural prior which handles order variations, background/redundant frames, and repeated actions in a single global alignment framework. In addition, we develop a unified optimal transport-based approach for joint video alignment and action segmentation.
@@
Action Segmentation. Initial works in self-supervised action segmentation perform representation learning and offline clustering as disjoint steps. Please see Ding et al. [16] for a recent survey. CTE [34] trains a temporal embedding first and then employs K-Means to cluster the embedded representations. To enhance CTE [34], VTE [73] and ASAL [43] introduce a visual embedding and an action embedding respectively. The above methods separate representation learning and offline clustering, prohibiting effective communications between the two modules. Recently, methods which jointly conduct representation learning and online clustering have been developed, e.g., UDE [65], TOT [35], and UFSA [71]. However, their performance may deteriorate in cases of order variations, unbalanced segmentation, and repeated actions. To overcome these limitations, ASOT [80] introduces a fused Gromov-Wasserstein optimal transport formulation with a structural prior. Here, we propose a similar optimal transport formulation for video alignment, outperforming previous methods on video alignment datasets, and extend it to a joint video alignment and action segmentation model, establishing the new state of the art on action segmentation benchmarks.
@@
Optimal Transport with Structured Data. Optimal transport underpins several computer vision and machine learning applications. A comprehensive review of optimal transport and its applications in machine learning is presented in Khamis et al. [30]. Optimal transport applications in computer vision include keypoint matching [47, 56], point set registration [58], object detection [14], object tracking [40], video alignment [45], and procedure learning [10]. For problems with structured data, a Gromov-Wasserstein optimal transport with a structural prior is frequently employed, e.g., graph matching [79], brain image registration [67], and action segmentation [80]. In this paper, we develop a Gromov-Wasserstein optimal transport with a structural prior for video alignment and extend it to joint video alignment and action segmentation. To our best knowledge, our work is the first to incorporate video alignment and action segmentation into a unified optimal transport framework.
@@ 3 Our Approach
Notations. First of all, $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}A_{ij}B_{ij}$ denotes the dot product of $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times m}$, $\mathbf{1}_{n}\in\mathbb{R}^{n}$ models a vector of ones, and $[n]=\{1,\dots,n\}$ represents a discrete set of $n$ elements. Next, $\Delta_{m}\subset\mathbb{R}^{m}$ denotes the $(m-1)$ dimensional probability simplex, while $\Delta_{m}^{n}\subset\mathbb{R}^{m\times n}$ models the Cartesian product space consisting of $n$ such simplexes. Furthermore, $X=\{x_{1},\dots,x_{N}\}$ and $Y=\{y_{1},\dots,y_{M}\}$ denote two input videos of $N$ and $M$ frames respectively. Let $f_{\bm{\theta}}$ with learnable parameters $\bm{\theta}$ be the embedding function, frame-level embeddings of $X$ and $Y$ are expressed as $\mathbf{X}=f_{\bm{\theta}}(X)\in\mathbb{R}^{N\times D}$ and $\mathbf{Y}=f_{\bm{\theta}}(Y)\in\mathbb{R}^{M\times D}$, where $D$ is the embedding vector length. Lastly, $K$ learnable action centroids are expressed by $\mathbf{A}=\left[\mathbf{a}_{1},\dots,\mathbf{a}_{K}\right]\in\mathbb{R}^{D\times K}$.
@@ 3.1.1 Optimal Transport with Structured Data
Kantorovich Optimal Transport. We briefly describe the conventional optimal transport formulation, also known as Kantorovich optiomal transport (KOT) [66], in the discrete setting. The KOT problem aims to find the minimum-cost coupling $\mathbf{T}^{\star}$ between histograms $\mathbf{p}\in\Delta_{n}$ and $\mathbf{q}\in\Delta_{m}$ with a ground cost $\mathbf{C}\in\mathbb{R}^{n\times m}_{+}$ and is written as:
@@
Gromov-Wasserstein Optimal Transport. For histograms defined over incomparable spaces, Gromov-Wasserstein (GW) optimal transport [52] is typically employed as:
@@
Fused Gromov-Wasserstein Optimal Transport. Fused Gromov-Wasserstein (FGW) optimal transport [68, 72] which merges KOT and GW formulations is often used for problems with known ground cost and structural prior. Let $\alpha\in[0,1]$, the FGW problem is expressed as:
@@
Balanced Optimal Transport. The above optimal transport problems impose balanced assignment constraints $\mathbf{T}\in\mathcal{T}_{\mathbf{p},\mathbf{q}}=\{\mathbf{T}\in\mathbb{R}_{+}^{n\times m}\mid\mathbf{T}\mathbf{1}_{m}=\mathbf{p},\mathbf{T}^{\top}\mathbf{1}_{n}=\mathbf{q}\}$. Recent works have relaxed these constraints by replacing (one [80] or both [67]) marginal constraints on $\mathbf{T}$ with penalty terms in the objective, yielding (partial [80] or full [67]) unbalanced assignment constraints. For video alignment, we adopt the full unbalanced formulation [67] but the results are worse than those of the balanced formulation, as we will show later in Sec. 4.1. This is likely because it is difficult to balance multiple losses (the full unbalanced formulation yields two extra penalty terms) and video alignment is generally more balanced than action segmentation (the number of frames is much larger than the number of actions).
@@ 3.1.2 Video Alignment Optimal Transport
Here, $\delta_{ik}=|i-k|$, $\delta_{jl}=|j-l|$, and a radius parameter $r\in(0,1]$. The GW component encourages temporal consistency over $\mathbf{T}$. In particular, assigning temporally nearby frames in $X$ ($\delta_{ik}\leq Nr$) to temporally distant frames in $Y$ ($\delta_{jl}>Mr$) incurs a cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=\frac{1}{r}$), whereas mapping temporally nearby frames in $X$ ($\delta_{ik}\leq Nr$) to temporally adjacent frames in $Y$ ($\delta_{jl}\leq Mr$) or mapping temporally distant frames in $X$ ($\delta_{ik}>Nr$) to temporally remote frames in $Y$ ($\delta_{jl}>Mr$) incurs no cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=0$). The GW component is capable of handling order variations and repeated actions, as shown in ASOT [80].
Fast Numerical Solver for VAOT. The GW component in Eq. 3 can be computed efficiently as $\mathcal{F}_{\text{GW}}(\mathbf{C}^{x},\mathbf{C}^{y},\mathbf{T})=\langle\mathbf{C}^{x}\mathbf{T}\mathbf{C}^{y},\mathbf{T}\rangle$ since the cost function $L(a,b)=ab$ can be factorized [52]. In addition, by adding an entropy regularization term $-\epsilon H(\mathbf{T})$, with $H(\mathbf{T})=-\sum_{i,j}T_{ij}\log T_{ij}$ and $\epsilon>0$, to the FGW formulation in Eq. 3, we can obtain the solution $\mathbf{T}^{\star}$ efficiently via projected mirror descent [52], which can be run on GPUs. Our solver often converges in less than $25$ iterations. By exploiting the sparse structures of $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$, each iteration has $O(NM)$ time complexity.
Background/Redundant Frame Handling. To tackle background/redundant frames, we follow VAVA [45] to add a virtual frame to $X$ and $Y$ so that background/redundant frames are explicitly assigned to it. Specifically, we append an extra row and column to $\mathbf{T}$ and expand other variables accordingly. If the assignment probability of $x_{i}$ ($i\leq N$) with every $y_{j}$ ($j\leq M$) is smaller than a threshold parameter $\zeta$, we match $x_{i}$ with the virtual frame $y_{M+1}$. Similarly, if the assignment probability of $y_{j}$ ($j\leq M$) with every $x_{i}$ ($i\leq N$) is smaller than $\zeta$, we match $y_{j}$ with $x_{N+1}$. Note that virtual frames and their associated frames are excluded from computing the losses. As shown in Sec. 4.1, handling background/redundant frames leads to performance gain.

### v3

@@ Abstract
We introduce a novel approach for simultaneous self-supervised video alignment and action segmentation based on a unified optimal transport framework. In particular, we first tackle self-supervised video alignment by developing a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which trains efficiently on GPUs and needs only a few iterations for solving the optimal transport problem. Our single-task method achieves the state-of-the-art performance on multiple video alignment benchmarks and outperforms VAVA, which relies on a traditional Kantorovich optimal transport formulation with an optimality prior. Furthermore, we extend our approach by proposing a unified optimal transport framework for joint self-supervised video alignment and action segmentation, which requires training and storing a single model and saves both time and memory consumption as compared to two different single-task models. Extensive evaluations on several video alignment and action segmentation datasets demonstrate that our multi-task method achieves comparable video alignment yet superior action segmentation results over previous methods in video alignment and action segmentation respectively. Finally, to the best of our knowledge, this is the first work to unify video alignment and action segmentation into a single model. Our code is available on our research website: https://retrocausal.ai/research/.
@@ 1 Introduction
Though the past decade has witnessed remarkable progress in human activity understanding in videos, the majority of the research efforts have been invested in action recognition [carreira2017quo, tran2018closer, wang2018non, feichtenhofer2019slowfast], which categorizes simple actions in short videos. In this paper, we study the two less-explored problems, i.e., temporal video alignment (frame-to-frame assignment), which establishes framewise correspondences between long videos recording a complex activity, and temporal action segmentation (frame-to-action assignment), which assigns frames of long videos capturing a multi-phase activity to phase/action labels. Since acquiring per-frame annotations for supervised training is generally difficult and costly, we are interested in self-supervised approaches for video alignment and action segmentation.
@@
One popular group of self-supervised video alignment methods rely on global alignment techniques widely used in time series literature. For example, LAV [haresh2021learning] utilizes dynamic time warping [cuturi2017soft] by assuming monotonic orderings and no background/redundant frames. VAVA [liu2022learning] relaxes the above assumptions by incorporating an optimality prior into a standard Kantorovich optimal transport framework [cuturi2013sinkhorn], along with an inter-video contrastive term and an intra-video contrastive term. However, it is challenging to balance multiple losses as well as handle repeated actions. Similarly, self-supervised action segmentation methods based on optimal transport have been introduced, including TOT [kumar2022unsupervised] and UFSA [tran2024permutation] which may suffer in cases of order variations, unbalanced segmentation, and repeated actions. ASOT [xu2024temporally] addresses these drawbacks via a fused Gromov-Wasserstein optimal transport framework with a structural prior, which outperforms previous works in self-supervised action segmentation. Lastly, though both self-supervised video align and action segmentation require fine-grained temporal understanding of videos, their interaction in a multi-task learning setup has not been explored.
@@
Motivated by the above observations, we first propose VAOT (see Fig. 1(a)), a novel self-supervised video alignment approach based on a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which tackles order variations, background/redundant frames, and repeated actions in a single global alignment framework. Our single-task model trains efficiently on GPUs and requires few iterations to derive the optimal transport solution, while outperforming previous methods, including VAVA [liu2022learning], on video alignment datasets. Moreover, we develop VASOT (see Fig. 1(b)), a joint self-supervised video alignment and action segmentation approach by exploring the relationship between self-supervised video alignment and action segmentation via a unified optimal transport framework. Our multi-task model performs on par with prior works on video alignment benchmarks yet establishes the new state of the art on action segmentation benchmarks. In addition, our joint model requires training and storing a single model and saves both time and memory consumption as compared to two separate single-task models. Lastly, we observe in Sec. 4.3 that, in a multi-task learning setting, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance significantly.
@@ 2 Related Work
Self-Supervised Learning. Early self-supervised learning methods focus on designing image-based pretext tasks with pseudo-labels as supervision signals for learning representations such as image colorization [larsson2016learning, larsson2017colorization], object counting [noroozi2017representation, liu2018leveraging], predicting rotations [gidaris2018unsupervised], solving puzzles [carlucci2019domain, kim2019self], image inpainting [jenni2020steering], and image clustering [caron2018deep, caron2019unsupervised]. The above image-based approaches mostly extract spatial cues from the image content. Recently, significant efforts have been invested in video-based self-supervised learning methods, which exploit both spatial and temporal information. Examples of video-based pretext tasks include forecasting future frames [srivastava2015unsupervised, vondrick2016generating, ahsan2018discrimnet, diba2019dynamonet], enforcing temporal coherence [mobahi2009deep, zou2011unsupervised, goroshin2015unsupervised], predicting temporal order [misra2016shuffle, lee2017unsupervised, fernando2017self, xu2019self], arrow of time [pickup2014seeing, wei2018learning], and pace [benaim2020speednet, wang2020self, yao2020video], and utilizing contrastive learning [feichtenhofer2021large, hu2021contrast, qian2021spatiotemporal, dave2022tclr]. More recently, skeleton-based self-supervised learning methods with skeleton-based pretext tasks, e.g., skeleton inpainting [zheng2018unsupervised], motion prediction [su2020predict], skeleton sequence alignment [kwon2022context, tran2024learning], and utilizing neighborhood consistency [si2020adversarial], motion continuity [su2021self], and multiple pretext tasks [lin2020ms2l], have been introduced. These skeleton-based approaches may suffer from human pose estimation errors and missing context details. Here, we leverage video alignment and/or action segmentation as our video-based pretext tasks.
@@
Video Alignment. Self-supervised video alignment has attracted a great amount of research interest in recent years. TCC [dwibedi2019tcc] enforces cycle consistencies between corresponding frames across videos for learning representations. Recently, GTCC [donahue2024learning] extends TCC [dwibedi2019tcc] by proposing multi-cycle consistencies for tackling repeated actions. Both TCC [dwibedi2019tcc] and GTCC [donahue2024learning] perform local alignment by aligning each frame separately. Motivated by global alignment techniques for time series, methods which align the video as a whole have been introduced. LAV [haresh2021learning] which assumes monotonic orderings and no background/redundant frames relies on dynamic time warping [cuturi2017soft]. To handle non-monotonic orderings and background/redundant frames, VAVA [liu2022learning] employs a traditional Kantorovich optimal transport formulation [cuturi2013sinkhorn] with an optimality prior. In this work, we propose a fused Gromov-Wasserstein optimal transport formulation with a structural prior which handles order variations, background/redundant frames, and repeated actions in a single global alignment framework. In addition, we develop a unified optimal transport-based approach for joint video alignment and action segmentation.
@@
Action Segmentation. Initial works in self-supervised action segmentation perform representation learning and offline clustering as disjoint steps. Please see Ding et al. [ding2023temporal] for a recent survey. CTE [kukleva2019unsupervised] trains a temporal embedding first and then employs K-Means to cluster the embedded representations. To enhance CTE [kukleva2019unsupervised], VTE [vidalmata2021joint] and ASAL [li2021action] introduce a visual embedding and an action embedding respectively. The above methods separate representation learning and offline clustering, prohibiting effective communications between the two modules. Recently, methods which jointly conduct representation learning and online clustering have been developed, e.g., UDE [swetha2021unsupervised], TOT [kumar2022unsupervised], and UFSA [tran2024permutation]. However, their performance may deteriorate in cases of order variations, unbalanced segmentation, and repeated actions. To overcome these limitations, ASOT [xu2024temporally] introduces a fused Gromov-Wasserstein optimal transport formulation with a structural prior. Here, we propose a similar optimal transport formulation for video alignment, outperforming previous methods on video alignment datasets, and extend it to a joint video alignment and action segmentation model, establishing the new state of the art on action segmentation benchmarks.
@@
Optimal Transport with Structured Data. Optimal transport underpins several computer vision and machine learning applications. A comprehensive review of optimal transport and its applications in machine learning is presented in Khamis et al. [khamis2024scalable]. Optimal transport applications in computer vision include keypoint matching [liu2020semantic, sarlin2020superglue], point set registration [shen2021accurate], object detection [de2023unbalanced], object tracking [lee2024sota], video alignment [liu2022learning], and procedure learning [chowdhuryopel]. For problems with structured data, a Gromov-Wasserstein optimal transport with a structural prior is frequently employed, e.g., graph matching [xu2019gromov], brain image registration [thual2022aligning], and action segmentation [xu2024temporally]. In this paper, we develop a Gromov-Wasserstein optimal transport with a structural prior for video alignment and extend it to joint video alignment and action segmentation. To our best knowledge, our work is the first to incorporate video alignment and action segmentation into a unified optimal transport framework.
@@ 3 Our Approach
Notations. First of all, $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}A_{ij}B_{ij}$ denotes the dot product of $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times m}$, $\mathbf{1}_{n}\in\mathbb{R}^{n}$ models a vector of ones, and $n]=\{1,\dots,n\}$ represents a discrete set of $n$ elements. Next, $\Delta_{m}\subset\mathbb{R}^{m}$ denotes the (m1)(m-1) dimensional probability simplex, while $\Delta_{m}^{n}\subset\mathbb{R}^{m\times n}$ models the Cartesian product space consisting of $n$ such simplexes. Furthermore, $X=\{x_{1},\dots,x_{N}\}$ and $Y=\{y_{1},\dots,y_{M}\}$ denote two input videos of $N$ and $M$ frames respectively. Let $f_{\bm{\theta}}$ with learnable parameters $\bm{\theta}$ be the embedding function, frame-level embeddings of $X$ and $Y$ are expressed as $\mathbf{X}=f_{\bm{\theta}}(X)\in\mathbb{R}^{N\times D}$ and $\mathbf{Y}=f_{\bm{\theta}}(Y)\in\mathbb{R}^{M\times D}$, where $D$ is the embedding vector length. Lastly, $K$ learnable action centroids are expressed by $\mathbf{A}=\left[\mathbf{a}_{1},\dots,\mathbf{a}_{K}\right]\in\mathbb{R}^{D\times K}$.
@@ 3.1.1 Optimal Transport with Structured Data
Kantorovich Optimal Transport. We briefly describe the conventional optimal transport formulation, also known as Kantorovich optiomal transport (KOT) [thorpe2018introduction], in the discrete setting. The KOT problem aims to find the minimum-cost coupling $\mathbf{T}^{\star}$ between histograms $\mathbf{p}\in\Delta_{n}$ and $\mathbf{q}\in\Delta_{m}$ with a ground cost $\mathbf{C}\in\mathbb{R}^{n\times m}_{+}$ and is written as:
@@
Gromov-Wasserstein Optimal Transport. For histograms defined over incomparable spaces, Gromov-Wasserstein (GW) optimal transport [peyre2016gromov] is typically employed as:
@@
Fused Gromov-Wasserstein Optimal Transport. Fused Gromov-Wasserstein (FGW) optimal transport [titouan2019optimal, vayer2020fused] which merges KOT and GW formulations is often used for problems with known ground cost and structural prior. Let $\alpha\in[0,1]$, the FGW problem is expressed as:
@@
Balanced Optimal Transport. The above optimal transport problems impose balanced assignment constraints $\mathbf{T}\in\mathcal{T}_{\mathbf{p},\mathbf{q}}=\{\mathbf{T}\in\mathbb{R}_{+}^{n\times m}\mid\mathbf{T}\mathbf{1}_{m}=\mathbf{p},\mathbf{T}^{\top}\mathbf{1}_{n}=\mathbf{q}\}$. Recent works have relaxed these constraints by replacing (one [xu2024temporally] or both [thual2022aligning]) marginal constraints on $\mathbf{T}$ with penalty terms in the objective, yielding (partial [xu2024temporally] or full [thual2022aligning]) unbalanced assignment constraints. For video alignment, we adopt the full unbalanced formulation [thual2022aligning] but the results are worse than those of the balanced formulation, as we will show later in Sec. 4.1. This is likely because it is difficult to balance multiple losses (the full unbalanced formulation yields two extra penalty terms) and video alignment is generally more balanced than action segmentation (the number of frames is much larger than the number of actions).
@@ 3.1.2 Video Alignment Optimal Transport
Here, $\delta_{ik}=|i-k|$, $\delta_{jl}=|j-l|$, and a radius parameter $r\in(0,1]$. The GW component encourages temporal consistency over $\mathbf{T}$. In particular, assigning temporally nearby frames in $X$ ($\delta_{ik}\leq Nr$) to temporally distant frames in $Y$ ($\delta_{jl}>Mr$) incurs a cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=\frac{1}{r}$), whereas mapping temporally nearby frames in $X$ ($\delta_{ik}\leq Nr$) to temporally adjacent frames in $Y$ ($\delta_{jl}\leq Mr$) or mapping temporally distant frames in $X$ ($\delta_{ik}>Nr$) to temporally remote frames in $Y$ ($\delta_{jl}>Mr$) incurs no cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=0$). The GW component is capable of handling order variations and repeated actions, as shown in ASOT [xu2024temporally].
Fast Numerical Solver for VAOT. The GW component in Eq. 3 can be computed efficiently as $\mathcal{F}_{\text{GW}}(\mathbf{C}^{x},\mathbf{C}^{y},\mathbf{T})=\langle\mathbf{C}^{x}\mathbf{T}\mathbf{C}^{y},\mathbf{T}\rangle$ since the cost function $L(a,b)=ab$ can be factorized [peyre2016gromov]. In addition, by adding an entropy regularization term $\epsilon H(\mathbf{T})$, with $H(\mathbf{T})=-\sum_{i,j}T_{ij}\log T_{ij}$ and $\epsilon>0$, to the FGW formulation in Eq. 3, we can obtain the solution $\mathbf{T}^{\star}$ efficiently via projected mirror descent [peyre2016gromov], which can be run on GPUs. Our solver often converges in less than $25$ iterations. By exploiting the sparse structures of $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$, each iteration has $O(NM)$ time complexity.
Background/Redundant Frame Handling. To tackle background/redundant frames, we follow VAVA [liu2022learning] to add a virtual frame to $X$ and $Y$ so that background/redundant frames are explicitly assigned to it. Specifically, we append an extra row and column to $\mathbf{T}$ and expand other variables accordingly. If the assignment probability of $x_{i}$ ($i\leq N$) with every $y_{j}$ ($j\leq M$) is smaller than a threshold parameter $\zeta$, we match $x_{i}$ with the virtual frame $y_{M+1}$. Similarly, if the assignment probability of $y_{j}$ ($j\leq M$) with every $x_{i}$ ($i\leq N$) is smaller than $\zeta$, we match $y_{j}$ with $x_{N+1}$. Note that virtual frames and their associated frames are excluded from computing the losses. As shown in Sec. 4.1, handling background/redundant frames leads to performance gain.